# OVA-Det: Open Vocabulary Aerial Object Detection with Image-Text Collaboration

### v2

@@ Abstract
Aerial object detection plays a crucial role in numerous applications. However, most existing methods focus on detecting predefined object categories, limiting their applicability in real-world open scenarios. In this paper, we extend aerial object detection to open scenarios through image-text collaboration and propose OVA-Det, a highly efficient open-vocabulary detector for aerial scenes. Specifically, we first introduce an image-to-text alignment loss to replace the conventional category regression loss, thereby eliminating category limitations. Next, we propose a lightweight text-guided strategy that enhances the feature extraction process in the encoder and enables queries to focus on class-relevant image features within the decoder, further improving detection accuracy without introducing significant additional costs. Extensive comparison experiments demonstrate that the proposed OVA-Det outperforms state-of-the-art methods on all three widely used benchmark datasets by a large margin. For instance, for zero-shot detection on DIOR, OVA-Det achieves 37.2 mAP and 79.8 Recall, 12.4 and 42.0 higher than that of YOLO-World. In addition, the inference speed of OVA-Det reaches 36 FPS on RTX 4090, meeting the real-time detection requirements for various applications. The code is available at https://github.com/GT-Wei/OVA-Det.
{NoHyper}footnotetext: This work was done during Guoting Wei’s internship at IntelliFusion.
1 Introduction
Figure 1: Comparison of OVA-Det with other advanced open-vocabulary detectors on the DIOR, DOTA, and xView datasets in terms of Recall, mAP, and FPS. Marker size encodes FPS, such that biger markers indicate faster inference. All models are evaluated in a zero-shot detection setting.
Aerial object detection, which involves localizing and categorizing objects of interest within aerial images, has become increasingly important due to its broad application requirements, including earth monitoring, disaster search, and rescue [21, 7, 30]. Most existing aerial detectors [38, 19, 15, 25] can only identify predefined categories from their training sets, which strictly restrict their deployment in open-world scenarios, which have become key considerations for real-world applications. A major obstacle for these detectors in discovering novel classes is their reliance on mapping image features to a fixed set of class indices[12].
@@ 1 Introduction
However, in the field of aerial object detection, such work remains very limited. Aerial object detection poses unique challenges, including predominantly small objects, varying viewpoints, and limited detection data. Directly applying existing open-vocabulary detection methods designed for conventional perspectives to aerial perspectives may result in suboptimal performance. Very recently, some researchers have attempted to address this gap and proposed insightful methods [23, 40, 26]. The introduction of these works has validated the potential of the open-vocabulary approach in the aerial object detection field, but there is still significant room for optimization. For instance, DescReg [40] introduces a description regularization method that focuses on enhancing textual descriptions but overlooks the importance of visual-semantic interaction. CastDet [23] adopts a complex multi-teacher framework based on a two-stage detector for open-vocabulary detection, which neglects the efficiency.
In this paper, we propose an efficient open-vocabulary detection model specifically designed for aerial photography scenes, termed OVA-Det. Specifically, we adopt the basic architecture of RT-DETR, due to RT-DETR has high inference efficiency and concise structure. First, we adapt the pretrained CLIP model and integrate it into the base framework to leverage its image-text alignment capability and powerful feature extraction, alleviating the limited sample problem. We then propose an image-to-text alignment loss to bridge the domain gap between natural images and aerial images, extending the model’s applicability from closed-world to open-world scenarios. Next, considering the challenges posed by long perception distances in aerial images, such as small detection targets with less prominent visual features and strong background interference, we propose to enhance visual feature and object queries in encoder and decoder with lightweight text guidance architectures, which further improving the detection accuracy without substantially increasing computational costs. Finally, extensive experiments are conducted to validate our method on three widely used benchmark datasets for zero-shot aerial object detection, as well as on two commonly adopted datasets for traditional closed-set aerial object detection. As illustrated in Figure 1, our proposed OVA-Det outperforms YOLO-World and DescReg in terms of mAP, Recall, and inference speed, demonstrating the effectiveness of our design.
@@
  2. We introduce lightweight text-guidance architectures, text-guided feature enhancement (TG-FE) and text-guided query enhancement (TG-QE) to enhance object visual features while suppressing background interference, further improving detection accuracy.
@@
  3. Our method demonstrates SOTA on multiple widely used benchmark datasets for zero-shot detection, as well as in traditional aerial object detection tasks, while maintaining a high inference speed. For instance, the proposed OVA-Det achieves 37.2 mAP and 79.8 Recall, while maintaining 36 FPS. To the best of our knowledge, we are the first real time open-vocabulary aerial detection method.
Figure 2: Overall architecture of OVA-Det. The proposed OVA-Det can be summarized into two main components: the Image-to-Text Alignment and the Text-Guided Strategy. (1) Image-to-Text Alignment connects visual objects with class embeddings rather than relying on class indices, which enables the model to detect classes beyond a predefined set. It involves a text encoder, a contrastive head, and a image-text alignment loss. (2) Text-Guided Strategy explores cross-modal interaction to enhance class-relevant feature extraction and suppress background interference, forming a text-guided encoder-decoder structure.
2 Related Work
2.1 Open Vocabulary Object Detection
Traditional object detectors [29, 1, 2, 46] predominantly rely on fully supervised training, mapping image features to class indices [12]. Such detectors are limited to recognizing only the categories seen during training, known as fixed-vocabulary (closed-set) detection. In contrast, open-vocabulary object detection (OVD) aims to explore the relationship between image and text to enable detectors to identify classes beyond those in the training sets.
@@ 2.1 Open Vocabulary Object Detection
Inspired by CLIP [28, 17] establish image-text feature alignment, recent works have incorporated class semantic information into detectors, exploring open-vocabulary object detection. These approaches can be roughly divided into two types. The first type [11, 45, 34, 20, 47, 3] uses networks like FPN to extract object proposals, which are then classified using pre-trained VLMs. For example, VILD [11] enhances OVD capability by distilling knowledge from pre-trained VLMs. Similarly, VL-PLM [45] leverages pre-trained VLMs to generate region-text pair labels for unlabeled images, these labels are subsequently used to train the detector, enabling recognition of novel object categories not present in the original training set. The second type integrates detection and grounding tasks [22, 44, 24, 39, 4], aiming to collect sufficient data to train specialized VLMs from scratch for detection. For instance, GLIP [22] unifies object detection and phrase grounding tasks to pre-train a detector for open-vocabulary detection. Grounding DINO [24] incorporates the pre-training strategy into transformer detection [43] based on phrase grounding.
@@
The most relevant work is YOLO-World [4], which extends YOLO to the open-vocabulary detection field through visual-language modeling. Firstly, YOLO-World and OVA-Det are proposed for different domains, natural images and aerial images. Secondly, the module design and loss function are also different. Unlike YOLO-World, which primarily uses text as guidance and focuses only on the feature extraction part, our proposed OVA-Det leverages text information to guide and enhance features in both the encoder and decoder with a concise architecture. This design is more effective in addressing the unique challenges of aerial images, such as low-quality and various viewpoints issues. Furthermore, the image-to-text alignment loss in OVA-Det replaces bidirectional alignment with unidirectional alignment, reducing optimization difficulty. These differences allow OVA-Det to achieve better performance in the aerial domain.
2.2 Aerial Object Detection
@@
Previous studies [8, 15, 19, 38] on aerial object detection have predominantly employed deep neural networks and coarse-to-fine architectures to enhance detection performance. However, these methods, functioning as closed-set detectors, are inherently limited in their ability to recognize unseen categories. Only a few works [40, 23, 26] explore open-vocabulary detection for aerial images. DescReg [40] argues that aerial images exhibit a weak semantic-visual correlation and utilizes a triplet loss to preserve the visual similarity structure within the classification space, thereby enhancing knowledge transfer from base to novel classes. CastDet [23] employs a semi-supervised training paradigm within a complex multi-teacher framework. It leverages a pre-trained vision-language model as a teacher to generate high-quality pseudo-labels for unlabeled images containing novel classes, thereby enabling the detection of object categories beyond those present in the training dataset. LAE-DINO [26] introduces a DINO-based detector that leverages dynamic vocabulary construction and visual-guided text prompt learning to explore class semantics, thereby enhancing OVD performance. These methods demonstrate the potential of the open-vocabulary approach in aerial object detection.
@@
However, existing methods rely on complex architectures that overlook the efficiency demands in aerial detection and fail to fully exploit the benefits of image-text collaboration. In response, OVA-Det builds upon a lighter framework and adopts a text-guided strategy to propagate class embeddings as clues throughout the entire network, helping the model focus on class-relevant object features and suppress extensive background interference.
3 Method
3.1 Model Architecture
@@
The overall architecture of OVA-Det is illustrated in Fig. 2. From the figure, we can see that the designs in the proposed OVA-Det focus on two major components: the image-to-text alignment component and the text-guided strategy component. The former is proposed to overcome category limitations. We integrate class embeddings via a text encoder, constructing both a contrastive head and an image-to-text alignment loss to facilitate image-text alignment. The latter is designed to enhance image features while suppressing background interference. We exploit image-text fusion between image features and class embeddings. The key components of the fusion modules are Text-Guided Feature Enhancement (TG-FE) and Text-Guided Query Enhancement (TG-QE), forming a text-guided encoder-decoder structure.
3.2 Image-Text Alignment
@@
3.2.1 Text Encoder
To incorporate class semantics into the detector, we adopt the CLIP text encoder [28, 5]. Since it contains very rich text features and is abundant for aerial image detection. The encoder accepts a class vocabulary $C\in\mathbb{R}^{n}$ and produces class embeddings $T=\text{TextEncoder}(C)\in\mathbb{R}^{n\times d}$, where $n$ represents the number of class names and $d$ denotes the dimensions of the class embeddings.
3.2.2 Image encoder
Unlike natural images, the field of aerial object detection lacks sufficient data to support training an open-vocabulary detector from scratch. To enhance visual-semantic integration and bridge the connection between visual and text features, we also incorporate the CLIP image encoder [28, 5]. Specifically, to ensure efficiency, we select the ResNet-50 version, modify it by removing the final AttnPool layer, and return the last three layers of the backbone as multi-scale feature outputs. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the extracted image features $\{F_{3},F_{4},F_{5}\}=\text{ImageEncoder}(I)\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$, where $H^{\prime},W^{\prime},$ and $C$ respectively denote the height, width, and channels of the feature maps.
3.2.3 Image-to-Text Alignment Loss
@@
where $\mathbf{v}$ is the query embedding, $\mathbf{t}$ is the text embedding, and $\mathrm{proj}(\cdot)$ denotes the linear layer used to align image features with class embeddings. $\alpha$ and $\beta$ are learnable scale and offset parameters, respectively.
This design allows us to fine-tune only a small number of learnable parameters, reducing optimization difficulty and the need for training samples, which is beneficial for overcoming the challenge of limited training samples in aerial imagery.
Detection Loss. The OVA-Det loss consists of two components: a image-to-text alignment loss and a location loss. Specifically, given an image and a set of class vocabularies, OVA-Det outputs $N$ object predictions $\{(b_{i},s_{i})\}_{i=1}^{N}$ for each decoder, where $N$ denotes the number of queries. The corresponding ground-truth labels for these predictions are denoted as $\{(b_{k},u_{k})\}_{k=1}^{K}$.
@@
$\small\mathrm{L_{con}}(\theta,\mathbf{u})=\begin{cases}-\theta\bigl{(}u\log(\theta)+(1-\mathbf{u})\log(1-\theta)\bigr{)},&\text{if }\mathbf{u}>0,\\ -\alpha\,\theta^{\gamma}\,\log\bigl{(}1-\theta\bigr{)},&\text{if }\mathbf{u}=0,\end{cases}$ (2)
@@
3.3 Text-Guided Strategy
By incorporating the image-text alignment approach, we have already extended detection capabilities beyond predefined categories. In this subsection, we further exploit text information to help the model focus on class-relevant features and enhance image-text alignment. Specifically, we propose a Text-Guided strategy, which uses textual information as guidance to reinforce aerial image features and suppress background interference. We design two key components: Text-Guided Feature Enhancement (TG-FE) and Text-Guided Query Enhancement (TG-QE). The former acts on the feature extraction process and the latter works on the query, achieving simultaneous enhancement of both the encoder and decoder.
Figure 3: The Text-Guided Encoder in OVA-Det. The TG-FE enhance features extraction by incorporating relevant class semantics into image features.
3.3.1 Text-Guided Feature Enhancement
Due to the long imaging distance in aerial imagery, the images often contain small, low-resolution targets whose visual features are not prominent, making detection relatively challenging. Additionally, the broad coverage of aerial images means that the background may contain elements that strongly interfere with the detection process. To address these two issues, we propose using text guidance to enhance visual representations and design TG-FE to operate on the encoder part.

### v3

# RT-OVAD: Real-Time Open-Vocabulary Aerial Object Detection via Image-Text Collaboration

@@ Abstract
Aerial object detection plays a crucial role in numerous applications. However, most existing methods focus on detecting predefined object categories, limiting their applicability in real-world open scenarios. In this paper, we extend aerial object detection to open scenarios through image-text collaboration and propose RT-OVAD, the first real-time open-vocabulary detector for aerial scenes. Specifically, we first introduce an image-to-text alignment loss to replace the conventional category regression loss, thereby eliminating category constraints. Next, we propose a lightweight image–text collaboration strategy comprising an image–text collaboration encoder and a text-guided decoder. The encoder simultaneously enhances visual features and refines textual embeddings, while the decoder guides object queries to focus on class-relevant image features. This design further improves detection accuracy without incurring significant computational overhead. Extensive experiments demonstrate that RT-OVAD consistently outperforms existing state-of-the-art methods across open-vocabulary, zero-shot, and traditional closed-set detection tasks. For instance, on the open-vocabulary aerial detection benchmarks DIOR, DOTA-v2.0, and LAE-80C, RT-OVAD achieves 87.7 AP50, 53.8 mAP, and 23.7 mAP, respectively, surpassing the previous state-of-the-art (LAE-DINO) by 2.2, 7.0, and 3.5 points. In addition, RT-OVAD achieves an inference speed of 34 FPS on an RTX 4090 GPU, approximately three times faster than LAE-DINO (10 FPS), meeting the real-time detection requirements of diverse applications. The code will be released at https://github.com/GT-Wei/RT-OVAD.
Index Terms:
Open vocabulary aerial object detection, image-text collaboration, text-guided strategy.
I Introduction
Aerial object detection, which involves localizing and categorizing objects of interest within aerial images, has become increasingly important due to its broad application requirements, including earth monitoring, disaster search, and rescue [1, 2, 3]. Most existing aerial detectors [4, 5, 6, 7, 8, 9, 10] can only identify predefined categories from their training sets, which strictly restrict their deployment in open-world scenarios, which have become key considerations for real-world applications. A major obstacle for these detectors in discovering novel classes is their reliance on mapping image features to a fixed set of class indices[11].
Figure 1: Comparison of RT-OVAD with other advanced open-vocabulary detectors on the DIOR, DOTA-v2.0, and LAE-80C datasets in terms of mAP/AP50, FPS, and parameter count. Marker size indicates the number of parameters, with larger markers representing models with more parameters. All models are evaluated under the open-vocabulary aerial detection setting.
@@ I Introduction
However, in the field of aerial object detection, such work remains limited. Aerial object detection poses unique challenges, including predominantly small objects, varying viewpoints, background interference, and limited detection data. Directly applying existing open-vocabulary detection methods designed for conventional perspectives to aerial perspectives may result in suboptimal performance. Very recently, some researchers have attempted to address this gap and proposed insightful methods [18, 19, 20, 21]. These works have demonstrated the potential of the open-vocabulary approach in the aerial object detection field, but there is still significant room for optimization. First, image-text alignment poses a significant challenge for open-vocabulary aerial detection. Existing methods often adopt alignment strategies developed for natural-image-based detectors, which may lead to overfitting due to the limited aerial detection data. Secondly, current open-vocabulary aerial detectors typically overlook visual-semantic interaction and, in particular, fail to address the background interference that arises during this process, which is especially prominent in aerial imagery. Finally, there remains a scarcity of efficient end-to-end open-vocabulary aerial detectors. Most existing methods [19, 20] utilize complex semi-supervised training strategies that heavily rely on pseudo-label generation for discovering novel classes. Although LAE-DINO [21] achieves an end-to-end framework based on Grounding DINO, it still falls short in achieving real-time inference efficiency.
In this paper, we propose a real-time open-vocabulary detection model specifically designed for aerial photography scenes, termed RT-OVAD. Specifically, we adopt the basic architecture of RT-DETR, due to its high inference efficiency and concise structure. First, we adapt the pretrained CLIP model and integrate it into the base framework to leverage its image-text alignment capability and powerful feature extraction, alleviating the limited sample problem. We then propose an image-to-text alignment loss to bridge the domain gap between natural images and aerial images, extending the model’s applicability from closed-world to open-world scenarios. Next, considering the challenges posed by long perception distances in aerial images, such as small detection targets with less prominent visual features, strong background interference, and intra-class instance scale variations, we propose lightweight image–text collaboration architectures consisting of an image–text collaboration encoder and a text-guided decoder. The encoder enhances both visual and textual representations, while the decoder guides object queries to focus on class-relevant features. This design further improves detection accuracy without incurring substantial computational costs. Finally, we conduct comprehensive experiments across three benchmark tasks: open-vocabulary detection, zero-shot detection, and traditional closed-set detection. As illustrated in Figure 1, RT-OVAD consistently surpasses mainstream open-vocabulary detectors (GLIP, Grounding DINO, YOLO-World) as well as the recent aerial-specific detector LAE-DINO in terms of both mAP and inference speed, demonstrating the effectiveness of the proposed approach.
@@
  2. We introduce lightweight image–text collaboration architectures that enhances both visual and textual representations while suppressing background interference, further improving detection accuracy. This strategy comprises three specific modules: Text-Guided Feature Enhancement (TG-FE) and Visual-Guided Text Refinement (VG-TR), which operate in the encoder, and Text-Guided Query Enhancement (TG-QE), which operates in the decoder.
@@
  3. Our method achieves SOTA performance across multiple benchmark datasets for open-vocabulary and zero-shot aerial detection, as well as traditional closed-set tasks, while maintaining 34 FPS. To the best of our knowledge, RT-OVAD is the first real-time open-vocabulary aerial object detection method.
II Related Work
II-A Open Vocabulary Object Detection
@@
Traditional object detectors [22, 23, 24, 25] predominantly rely on fully supervised training, mapping image features to class indices [11]. Such detectors are limited to recognizing only the categories seen during training, known as fixed-vocabulary (closed-set) detection. In contrast, open-vocabulary object detection (OVD) aims to explore the relationship between image and text to enable detectors to identify classes beyond those in the training sets. Inspired by CLIP [12] establish image-text feature alignment, recent works have incorporated class semantic information into detectors, exploring open-vocabulary object detection. These approaches can be roughly divided into two types. The first type [26, 13, 14, 27, 28, 29, 30] uses networks like FPN to extract object proposals, which are then classified using pre-trained VLMs. For example, VILD [13] enhances OVD capability by distilling knowledge from pre-trained VLMs. Similarly, VL-PLM [14] leverages pre-trained VLMs to generate region-text pair labels for unlabeled images, these labels are subsequently used to train the detector, enabling recognition of novel object categories not present in the original training set. The second type integrates detection and grounding tasks [17, 31, 32, 15, 33, 16], aiming to collect sufficient data to train specialized VLMs from scratch for detection. For instance, GLIP [17] unifies object detection and phrase grounding tasks to pre-train a detector for open-vocabulary detection. Grounding DINO [15] incorporates the pre-training strategy into transformer detection [34] based on phrase grounding.
@@
The most relevant work is YOLO-World [16], which extends YOLO to the open-vocabulary detection field through visual-language modeling. Firstly, YOLO-World and RT-OVAD are proposed for different domains, natural images and aerial images. Secondly, the module design and loss function are also different. Unlike YOLO-World, which primarily uses text as guidance and focuses only on the feature extraction part, our proposed RT-OVAD leverages text information to guide and enhance features in both the encoder and decoder with a concise architecture. This design is more effective in addressing the unique challenges of aerial images, such as low-quality and various viewpoints issues. Furthermore, the image-to-text alignment loss in RT-OVAD replaces bidirectional alignment with unidirectional alignment, reducing optimization difficulty. These differences allow RT-OVAD to achieve better performance in the aerial domain.
II-B Aerial Object Detection
@@
Previous studies [4, 7, 6, 5, 8, 9] on aerial object detection have predominantly employed deep neural networks and coarse-to-fine architectures to enhance detection performance. However, these methods, functioning as closed-set detectors, are inherently limited in their ability to recognize unseen categories. Only a few works [18, 19, 20, 21] explore open-vocabulary detection for aerial images. DescReg [18] argues that aerial images exhibit a weak semantic-visual correlation and utilizes a triplet loss to preserve the visual similarity structure within the classification space, thereby enhancing knowledge transfer from base to novel classes. CastDet [19] employs a semi-supervised training paradigm within a complex multi-teacher framework. It leverages a pre-trained vision-language model as a teacher to generate high-quality pseudo-labels for unlabeled images containing novel classes, thereby enabling the detection of object categories beyond those present in the training dataset. LAE-DINO [21] utilizes vision-language models to expand detectable categories and introduces a DINO-based detector that leverages dynamic vocabulary construction and visual-guided text prompt learning to explore class semantics, thereby enhancing OVD performance.
@@
However, existing methods rely on complex architectures that overlook the efficiency demands in aerial detection and fail to fully exploit the benefits of image-text collaboration. In response, RT-OVAD builds upon a lighter framework and adopts a text-guided strategy to propagate class embeddings as clues throughout the entire network, helping the model focus on class-relevant object features and suppress extensive background interference.
Figure 2: Overall architecture of RT-OVAD. The proposed RT-OVAD can be summarized into two main components: the Image-to-Text Alignment and the Image-Text collaboration Strategy. (1) Image-to-Text Alignment connects visual objects with class embeddings rather than relying on class indices, which enables the model to detect classes beyond a predefined set. It involves a text encoder, a contrastive head, and a image-text alignment loss. (2) Image-Text collaboration Strategy explores cross-modal interaction to enhance class-relevant feature extraction, suppress background interference and alleviate the impact of intra-class scale variations, forming a image-text collaboration encoder-decoder structure.
III Method
III-A Model Architecture
@@
The overall architecture of RT-OVAD is illustrated in Fig. 2. From the figure, we can see that the designs in the proposed RT-OVAD focus on two major components: the image-to-text alignment component and the image-text collaboration component. The former is proposed to overcome category limitations. We integrate class embeddings via a text encoder, constructing both a contrastive head and an image-to-text alignment loss to facilitate image-text alignment. The latter is designed to enhance visual features, suppress background interference, and mitigate intra-class scale variations in aerial imagery. Specifically, we exploit image–text fusion between image features and class embeddings, forming an image–text collaborative encoder–decoder structure. This component consists of three key modules: Text-Guided Feature Enhancement (TG-FE), Visual-Guided Text Refinement (VG-TR), and Text-Guided Query Enhancement (TG-QE).
III-B Image-Text Alignment
@@
III-B1 Text Encoder
To incorporate class semantics into the detector, we adopt the CLIP text encoder [12]. Since it contains very rich text features and is abundant for aerial image detection. The encoder accepts a class vocabulary $C\in\mathbb{R}^{n}$ and produces class embeddings $T=\text{TextEncoder}(C)\in\mathbb{R}^{n\times d}$, where $n$ represents the number of class names and $d$ denotes the dimensions of the class embeddings.
III-B2 Image encoder
Unlike natural images, the field of aerial object detection lacks sufficient data to support training an open-vocabulary detector from scratch. To enhance visual-semantic integration and bridge the connection between visual and text features, we also incorporate the CLIP image encoder [12]. Specifically, to ensure efficiency, we select the ResNet-50 version, modify it by removing the final AttnPool layer, and return the last three layers of the backbone as multi-scale feature outputs. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the extracted image features $\{F_{3},F_{4},F_{5}\}=\text{ImageEncoder}(I)\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$, where $H^{\prime},W^{\prime},$ and $C$ respectively denote the height, width, and channels of the feature maps.
III-B3 Image-to-Text Alignment Loss
@@
where $\mathbf{v}$ is the query embedding, $\mathbf{t}$ is the text embedding enhanced by VG-TR (4), and $\mathrm{proj}(\cdot)$ denotes the linear layer used to align image features with class embeddings. $\alpha$ and $\beta$ are learnable scale and offset parameters, respectively.
This design allows us to fine-tune only a small number of learnable parameters, reducing optimization difficulty and the risk of overfitting, which is beneficial for overcoming the challenge of limited training samples in aerial imagery.
Detection Loss. The RT-OVAD loss consists of two components: a image-to-text alignment loss and a location loss. Specifically, given an image and a set of class vocabularies, RT-OVAD outputs $N$ object predictions $\{(b_{i},s_{i})\}_{i=1}^{N}$ for each decoder, where $N$ denotes the number of queries. The corresponding ground-truth labels for these predictions are denoted as $\{(b_{k},u_{k})\}_{k=1}^{K}$.
@@
$\small\mathrm{L_{con}}(\theta,\mathbf{u})=\begin{cases}-\mathbf{u}\bigl{(}\mathbf{u}\log(\theta)+(1-\mathbf{u})\log(1-\theta)\bigr{)},&\text{if }\mathbf{u}>0,\\ -\alpha\,\theta^{\gamma}\,\log\bigl{(}1-\theta\bigr{)},&\text{if }\mathbf{u}=0,\end{cases}$ (2)
@@
III-C Image-Text collaboration Strategy
By incorporating the image-text alignment approach, we have already extended detection capabilities beyond predefined categories. In this subsection, we further exploit image-text collaboration strategy to enhance image-text alignment and help the model focus on class-relevant features. Specifically, we propose Text-Guided Feature Enhancement (TG-FE) and Visual-Guided Text Refinement (VG-TR) within the encoder to simultaneously enhance visual features and refine textual embeddings. Additionally, we introduce a Text-Guided Query Enhancement (TG-QE) module within the decoder to enhance object queries, forming a complementary image–text collaboration encoder-decoder architecture.
III-C1 Image-Text Collaboration Encoder
Text-Guided Feature Enhancement. Due to the long imaging distance in aerial imagery, the images often contain small, low-resolution targets whose visual features are not prominent, making detection relatively challenging. Additionally, the broad coverage of aerial images means that the background may contain elements that strongly interfere with the detection process. To address these two issues, we propose using text guidance to enhance visual representations and design TG-FE to operate on the encoder part.
Figure 3: The Image-Text collaboration Encoder in RT-OVAD. The TG-FE module enhances feature extraction by incorporating relevant class semantics into the visual features. The VG-TR module further refines the textual representations by fusing the enhanced visual features from TG-FE into the corresponding class embeddings.